# The Connection between Dusty Star-Forming Galaxies and the First Massive Quenched Galaxies

### v1

@@ 4.3 Limitations and caveats
One likely cause of the underprediction of the high-redshift MQ population is the simplified AGN feedback and black hole growth prescription adopted in the Henriques et al. (2020) version of L-Galaxies, originally introduced by Croton et al. (2006). As shown in this study, AGN feedback is the primary mechanism for quenching star formation in both DSFGs and MQs at high-redshift. In the Croton et al. (2006) framework, SMBH growth occurs mainly via mergers, and AGN feedback operates solely on the hot gas halo, suppressing cooling flows without affecting the cold gas content of the galaxy directly. A more sophisticated treatment of black hole accretion and AGN-driven feedback—such as one capable of heating or ejecting cold gas, as in the Izquierdo-Villalba et al. (2022) version of L-Galaxies and the Lagos et al. (2024) version of SHARK SAM—may therefore resolve the strong discrepancy with observations.
@@ Acknowledgements.
PA-A thanks the Coordenaçõ de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES), for supporting his PhD scholarship (project 88882.332909/2020-01). RKC is grateful for support from the Leverhulme Trust via the Leverhulme Early Career Fellowship. The model calibration was run on computing cluster facilities provided by the Flatiron Institute. The Flatiron Institute is supported by the Simons Foundation. LSJ acknowledges the support from CNPq (308994/2021-3) and FAPESP (2011/51680-6). MCV acknowledges the FAPESP for supporting his PhD (2021/06590-0). PA-A and BG acknowledge support from the Independent Research Fund Denmark (DFF) under grant 4251-00086B. FV acknowledges support from the DFF under grant 3120-00043B.
Appendix A Timescales between the peak of sub-millimetre emission and subsequent quenching

### v2

@@ 4.3 Limitations and caveats
One likely cause of the underprediction of the high-redshift MQ population is the simplified AGN-feedback and black hole growth prescription adopted in the Henriques et al. (2020) version of L-Galaxies, originally introduced by Croton et al. (2006). In the Croton et al. (2006) framework, SMBH growth occurs mainly through mergers, which can lead to an underestimation of SMBH masses in the early Universe ($z\gtrsim 4$). In the context of the ‘no HIMF’ recalibrated version presented in Araya-Araya et al. (2025), a $\lesssim 3\%$ fraction (scaled by the mass ratio) of the cold gas is accreted during mergers, independently of the halo mass. Among all recalibrated variants, the ‘no HIMF’ model is, in fact, one of the most efficient at accreting cold gas onto SMBHs. However, because this efficiency is independent of halo mass, the resulting $M_{\rm BH}-M_{\star}$ relation remains almost constant at high redshift (see Figure 12), leading to an underprediction of SMBH masses in massive galaxies compared with other models (see Figure 13 Lagos et al., 2025) and observational data from Decarli et al. (2010a, b), Suh et al. (2020), and Poitevineau et al. (2023) at $z\sim 3$. A more sophisticated treatment of SMBH growth and formation, as implemented in the Izquierdo-Villalba et al. (2022) and Spinoso et al. (2023) versions of L-Galaxies, improves consistency with current observations. Nevertheless, our model performs reasonably well in reproducing the observed SMBH mass function and the $M_{\rm BH}-M_{\star}$ relation at low redshift.
As shown in this study, AGN feedback is the primary mechanism responsible for quenching star formation in both DSFGs and MQs at high redshift. In the Croton et al. (2006) model, AGN feedback operates exclusively on the hot gas halo, suppressing cooling flows without directly affecting the cold gas reservoir of the galaxy. In the ‘no HIMF’ recalibrated model, AGN feedback is stronger than in the default Henriques et al. (2020) version. However, it still underpredicts the fraction of quiescent galaxies at high redshift—likely by the underestimation of SMBH masses at early times, as discussed above. As shown by Lagos et al. (2024) and De Lucia et al. (2024) for the SHARK and GAEA [^5] In fact, the SMBH–related processes prescription adopted in GAEA builds on the Croton et al. (2006) framework (Fontanot et al., 2020), but additionally includes cold gas accretion as a channel for SMBH growth, as well as AGN-driven winds that directly affect the cold gas in galaxy disks. SAMs, implementing a more comprehensive model for AGN-driven feedback, capable, for example, of heating or ejecting cold gas, may help resolve the strong discrepancy with observed MQ abundances. Increasing the number density of MQs at high redshift would naturally enhance the predicted transition rate between DSFGs and MQs.
@@ Acknowledgements.
We thank the anonymous referee for constructive comments that helped improve the clarity of this work. PA-A thanks the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES), for supporting his PhD scholarship (project 88882.332909/2020-01). RKC is grateful for support from the Leverhulme Trust via the Leverhulme Early Career Fellowship. The model calibration was run on computing cluster facilities provided by the Flatiron Institute. The Flatiron Institute is supported by the Simons Foundation. LSJ acknowledges the support from CNPq (308994/2021-3) and FAPESP (2011/51680-6). MCV acknowledges the FAPESP for supporting his PhD (2021/06590-0). PA-A and BG acknowledge support from the Independent Research Fund Denmark (DFF) under grant 4251-00086B. FV acknowledges support from the DFF under grant 3120-00043B.
Appendix A Energy input due to feedback
The energy inputs due to SN and AGN feedback, $\Delta E_{\rm SN}$ and $\Delta E_{\rm AGN}$, (see Figure 8), are not direct outputs from L-Galaxies; however, they can be easily derived from other properties. To calculate the energy input from SN feedback, we use the outputs ReheatingRate ($\Delta M_{\rm reheat}$) and EjectionRate ($\Delta M_{\rm eject}$), which correspond to the mass rates of reheated and ejected gas, respectively. It is important to note that, in L-Galaxies, ejection occurs only if there is remaining energy after reheating. The total energy input is then given by Equation 3 — Adapted from Equation S23 in the Supplementary Material of Henriques et al. (2020):
$\Delta E_{\rm SN}=\frac{1}{2}(\Delta M_{\rm reheat}+\Delta M_{\rm eject})\,V_{200c}^{2}\ ,$ (3)
where $V_{200c}$ is the halo virial velocity.
The AGN feedback in the Henriques et al. (2020) version of L-Galaxies is based on the Croton et al. (2006) model. To obtain the energy input due to AGN feedback, we first estimate the accretion using Equation 4:
$\Delta M_{\rm BH}=k_{\rm AGN}\left(\frac{M_{\rm hot}}{10^{11}\rm{M_{\odot}}}\right)\left(\frac{M_{\rm BH}}{10^{8}\rm{M_{\odot}}}\right)\frac{H_{0}}{H(z)}\ ,$ (4)
where $k_{\rm AGN}$ is a free parameter ($k_{\rm AGN}=0.07\,\rm{M_{\odot}}{\rm yr}^{-1}$ in the recalibrated model used in this work), $M_{\rm hot}$ is the mass of hot gas, and $M_{\rm BH}$ is the supermassive black hole mass. Finally, the energy input due to AGN feedback is given by $\Delta E_{\rm AGN}=\eta\,\Delta M_{\rm BH}\,c^{2}$ (Equation S29 in the Supplementary Material of Henriques et al. (2020)), where $\eta=0.1$ and $c$ is the speed of light.
Appendix B Timescales between the peak of sub-millimetre emission and subsequent quenching
@@
Appendix C $_{\rm BH}-MM_{\star}$ Relation
For completeness and as a sanity check, we present in Figure 12 the $_{\rm BH}MM_{\star}$ relation for the ‘no HIMF’ model of Araya-Araya et al. (2025) at $z=0$ and $z=3$. At $z=0$, we compare our predictions with the observational relations from Reines and Volonteri (2015) and Greene et al. (2020). The observational estimates from Decarli et al. (2010a, b); Suh et al. (2020) and Poitevineau et al. (2023) at $2.0<z<3.5$ are compared against our predictions at $z=3$.
Figure 12: The $M_{\rm BH}-M_{\star}$ relation at $z=0$ (left) and 3 (right) as predicted by the ‘no HIMF’ recalibrated model (Araya-Araya et al., 2025). The light-blue areas and lines represent the $16^{\rm th}$, $50^{\rm th}$, and $84^{\rm th}$ percentiles of all galaxies in the simulated box. The black and red dashed lines indicates the $M_{\rm BH}-M_{\star}$ relations derived from observations by Reines and Volonteri (2015) and Greene et al. (2020), respectively. The predictions at $z=3$ are compared to observational estimates from Decarli et al. (2010a, b) (black dots), Suh et al. (2020) (gray squares) and Poitevineau et al. (2023) (purple diamonds).